# XQC: Well-conditioned Optimization Accelerates Deep Reinforcement Learning

### v1

@@ Abstract
Sample efficiency is a central property of effective deep reinforcement learning algorithms. Recent work has improved this through added complexity, such as larger models, exotic network architectures, and more complex algorithms, which are typically motivated purely by empirical performance. We take a more principled approach by focusing on the optimization landscape of the critic network. Using the eigenspectrum and condition number of the critic’s Hessian, we systematically investigate the impact of common architectural design decisions on training dynamics. Our analysis reveals that a novel combination of batch normalization (bn), weight normalization (wn), and a distributional cross-entropy (ce) loss produces condition numbers orders of magnitude smaller than baselines. This combination also naturally bounds gradient norms, a property critical for maintaining a stable effective learning rate under non-stationary targets and bootstrapping. Based on these insights, we introduce xqc: a well-motivated, sample-efficient deep actor-critic algorithm built upon soft actor-critic that embodies these optimization-aware principles. We achieve state-of-the-art sample efficiency across 55 proprioception and 15 vision-based continuous control tasks, all while using significantly fewer parameters than competing methods.
@@ 1 Introduction
Sample efficiency remains a major challenge in deep rl. Methods that can learn effectively from limited interactions are crucial for applying rl in domains such as robotics, where generating data on real hardware is costly and time-consuming. Recent advances in model-free rl have primarily been driven by a paradigm of scaling—larger networks, higher update-to-date ratio (utd) ratios, and ever-increasing computational budgets (Nikishin et al., 2022; D’Oro et al., 2022; Nauman et al., 2024; 2025; Lee et al., 2025a; b; Palenicek et al., 2025). These works have viewed architectural improvements primarily as a means to scale up stably and in turn improve sample efficiency. While proven effective, this ‘bigger is better’ paradigm comes at the cost of computational efficiency and often overlooks a more fundamental question: Can we improve performance not by adding complexity, but by creating a better-conditioned optimization problem?
To develop this principled understanding, we conduct a systematic investigation into three commonly used architectural components whose roles in rl are often guided by heuristics. First, we examine normalization layers. The rl community has converged mainly on layer normalisation (ln(Ba et al., 2016), primarily due to concerns about the batch dependency of bn (Ioffe & Szegedy, 2015), which until recently was thought to be problematic in the rl setting. Second, we consider wn (Lyle et al., 2024; Loshchilov et al., 2025; Palenicek et al., 2025) by periodically projecting the network’s weights to the unit sphere, permitted through the normalization layers’ scale invariance property. A technique known to improve the effective learning rate (elr(Van Laarhoven, 2017). Lastly, we study the critic’s loss function. Distributional critics using a ce loss have grown in popularity (Bellemare et al., 2017). The conventional argument for their adoption is that modeling the full distribution of returns provides a better learning signal compared to regression with a mean squared error (mse) loss (Farebrother et al., 2024); there is evidence this loss is easier to optimize (Imani & White, 2018).
@@ 2 Preliminaries
where $Q$ and $V$ are the ‘soft’ parametric critic and value functions, respectively. Minimizing this Bellman error effectively is key to the success of actor-critic methods (Sutton & Barto, 2018). This error is typically minimized with the mean squared regression loss and gradient-based optimization. The distributional c51 algorithm (Bellemare et al., 2017) reformulates the task as a classification problem. Instead of a scalar estimate, the function approximator outputs the logits of a categorical distribution over the full support of $Q$. The Bellmann error can then be minimized using a categorical ce loss, which has been shown to improve performance and stability (Farebrother et al., 2024).
@@
${\mathcal{L}}({\bm{\theta}}+\delta{\bm{\theta}})\approx{\mathcal{L}}({\bm{\theta}})+\nabla_{\bm{\theta}}{\mathcal{L}}({\bm{\theta}})\delta{\bm{\theta}}+\frac{1}{2}{\delta{\bm{\theta}}}^{\top}\nabla^{2}_{{\bm{\theta}}}{\mathcal{L}}({\bm{\theta}})\delta{\bm{\theta}}$ (2)
@@ 3.2 Why does cross-entropy outperform the squared error?
For proofs see Appendix I. In practice, we do not require weight decay to attain good performance, and as a result, the Hessian was observed to not always be positive definite. Nonetheless, these results provide a formal intuition for our empirical result that ce losses consistently report smaller condition numbers. This analysis directly motivates the design of our xqc algorithm, presented next.
@@ 5 Experiments
Figure 6: xqc achieves state-of-the-art sample efficiency across 55 propioceptive continuous control tasks. We report the iqm auc of normalized returns. Error bars denote 90% sbcis. The right column shows total aggregated performance across the benchmarks (55 tasks). xqc matches or outperforms strong baselines, especially on the hardest DMC and HB tasks, while using a simpler and smaller architecture (see Section 5.3).
@@ 5.1 Sample efficiency results
Reinforcement learning from proprioception. Figure 6 shows that xqc matches our outperforms strong baselines simba-v2, mrq, bro and crossq+wn on all 4 benchmarks. The rightmost column shows that on average xqc performs as well simba-v2 while using significantly less network parameters and a substantially simpler architecture Figure 5. Notably xqc shows exceptional performance on the most complex tasks HB and DMC-hard. These environments are known to induce notoriously difficult and ill-conditioned optimization landscapes. xqc’s superior performance and learning speed suggest that its well-conditioned critic—characterized by a stable elr and bounded gradients as shown in Section 5.2—is fundamentally better equipped to handle the non-stationary targets and bootstrapping errors inherent in these challenging domains.
@@ 5.3 Parameter and compute effciency
Figure 9: xqc is significantly more parameter and compute efficient & scales stably with utd, network depths, and widths. Left two columns: xqc is significantly more parameter- and compute-efficient than the competing baselines. $\sim 4.5\times$ fewer parameters and $\sim 5\times$ fewer flop/s than simba-v2 and bro and $100\times$ fewer flop/s than brc. Columns 3-5: xqc’s scaling in terms of utd, layer width, and layer depth. Increasing compute and model capacity generally improves or maintains performance, demonstrating that our well-conditioned design is robust to scaling. Results are aggregated across all 55 proprioception tasks and 5 seeds each. Right column: We compare the full xqc algorithm against three variants: one replacing bn with ln (xqc ln), one replacing the ce loss with an mse loss (xqc mse), and one without wn (xqc w/o wn). Each component’s removal results in a significant performance drop, showing their synergistic contribution.
xqc achives its competitive sample efficiency while requiring $\sim 4.5\times$ fewer parameters than simba-v2 (Figure 9). This parameter efficiency directly results in high computational efficiency with $\sim 5\times$ fewer flop/s than simba-v2 and bro and $100\times$ fewer flop/s than brc. We conjecture that xqc’s superior computational efficiency is rooted in its well-conditioned architecture.
@@ 5.4 XQC scaling behaviour and architecture ablations
Figure 9 colums 3–5 demonstrate xqc’s scaling ability. xqc improves performance with increasing utd at a similar slope to simba-v2, with a faction of the parameters (Section 5.3). Similarly, xqc improves or maintains performance with larger and deeper networks. These results demonstrate robustness towards different hyperparameters and xqc’s ability to scale stably, enabled by its well-conditioned architectural design. This property is desirable since increasing compute always represents a trade-off between sample efficiency and wall-clock or energy. It allows practitioners to use their available budget most efficiently. To demonstrate the synergy, we ablate each of the three main components of the proposed xqc architecture: bn, wn, and the ce distributional critic. Results for all 55 and 5 seeds are shown in Figure 9 with per benchmark ablations in Figure 16. Our analysis confirms that each of the components is vital, especially in the most difficult DMC-hard and HB suits. The largest drop in performance occurs when switching bn for ln. While using a mse critic loss has the second most significant influence, the removal of wn is still substantial, but shows the lowest overall impact. In summary, combining each of the three components is vital for xqc’s performance.
@@ 6 Related Work
Our work is positioned at the intersection of several key research areas in deep rl that all attempt to improve sample efficiency. A prevailing trend for improving sample efficiency has been scaling along different axes. Compute scaling, particularly increasing the utd ratio, has recently been a major focus. However, naively increasing utd can lead to instability and overfitting on early experience (Nikishin et al., 2022). Researchers have suggested many different regularizers to stably increase the utd; From full parameter-resets (Nikishin et al., 2022; D’Oro et al., 2022), to critic ensembles (Chen et al., 2021) and drop-out (Hiraoka et al., 2021), to normalization layers like ln (Hiraoka et al., 2021; Lyle et al., 2024; Nauman et al., 2024; Lee et al., 2025a), bn (Bhatt et al., 2024), hyper-spherical normalization (Hussing et al., 2024; Lee et al., 2025b) and spectral normalization (Bjorck et al., 2021). Recently, works have found that the combination of normalization layers together with wn can stabilize the elr, helping against loss of plasticity (Lyle et al., 2024) and also enable scaling rl to high utd ratios (Palenicek et al., 2025). Recent works have worked on developing scaling laws for rl (Rybkin et al., 2025; Fu et al., 2025). Network scaling, is another path authors are exploring to increase sample effieincy. Bhatt et al. (2024) showed that bn allowed them to significantly scale the layer width. Since then, authors have looked into specific ln-based architectures (Nauman et al., 2024; Lee et al., 2025a) and network sparsity (Ma et al., 2025). Lee et al. (2025a) propose a ‘simplicity bias’ score, computed using an FFT and scoring ’simplicity’ higher for functions with lower frequency content across random initializations. This score has no theoretical justification relating it to sample efficiency or generalization. Another line of research attempts to scale network sizes dynamically during training (Liu et al., 2025; Kang et al., 2025). Concurrent work Castanyer et al. (2025) combine second-order optimization and multi-skip residual connections to improve scaling and monitor the trace of the Hessian for deep value-based rl. Model-based methods scale computation by learning a separate dynamics model, which is then leveraged in the rl loop (Janner et al., 2019; Hafner et al., 2020; Hansen et al., 2024).
@@ 7 Conclusion & Future Work
In this work, we shifted the focus from the prevailing pure scaling goal in deep rl and instead focus on improving the critic’s optimization landscape. Through an eigenvalue analysis of the critic’s Hessian, we demonstrate that specific architectural choices, namely batch normalization, weight normalization, and a distributional cross-entropy loss, create a better optimization landscape with a condition number orders of magnitude smaller during learning. This superior conditioning translates directly into learning performance gains. We propose xqc, an algorithm embodying these principles, which achieves state-of-the-art sample efficiency across 70 continuous control tasks from proprioception and vision domains. xqc accomplishes this performance with significantly fewer parameters than competing methods, underscoring that a principled focus on optimization fundamentals can yield greater performance and efficiency than brute-force scaling alone.
@@ Acknowledgements
This research was funded by the research cluster “Third Wave of AI”, funded by the excellence program of the Hessian Ministry of Higher Education, Science, Research and the Arts, hessian.AI. This work was also supported by a UKRI/EPSRC Programme Grant [EP/V000748/1].
@@ Appendix A Hyperparameters
Table 1 summarizes all proprioception-based experiments’ hyperparameters. The heuristic discount factor is determined using the heuristic proposed by (Hansen et al., 2024). Specifically, it is computed as:
@@
Block design
Dense(dim)
BN
ReLU
Dense(dim)
ReLU
BN
Dense(dim)
ReLU
Dense(4 $\times$ dim)
Scaler
ReLU
Dense(dim)
L2 Norm
LERP
L2 Norm
Dense(dim)
LN
ReLU
Dense
LN
Skip Connection
Dense(dim)
LN
ReLU
Dense
LN
Skip Connection
Dense(dim)
LN
Activation
@@ 10 BRO
Action repeat
MuJoCo: 1
DMC: 2
HB: 2
Myo: 2
MuJoCo: 1
DMC: 2
HB: 2
Myo: 2
MuJoCo: 1
DMC: 2
HB: 2
Myo: 2
MuJoCo: 1
DMC: 2
HB: 2
Myo: 2
MuJoCo: 1
DMC: 1
HB: 1
Myo: 1
MuJoCo: 1
DMC: 2
HB: 2
Myo: 2
MuJoCo: 1
DMC: 2
HB: 2
Myo: 2
@@
Block design
Dense(dim)
BN
ReLU
Dense(dim)
ReLU
@@
Replay buffer capacity 10M 10M
@@
easy: linear(1.0, 0.1, 100K)
medium: linear(1.0, 0.1, 500K)
hard: linear(1.0, 0.1, 2M)
easy: linear(1.0, 0.1, 100K)
medium: linear(1.0, 0.1, 500K)
hard: linear(1.0, 0.1, 2M)

### v2

@@ Abstract
Sample efficiency is a central property of effective deep reinforcement learning algorithms. Recent work has improved this through added complexity, such as larger models, exotic network architectures, and more complex algorithms, which are typically motivated purely by empirical performance. We take a more principled approach by focusing on the optimization landscape of the critic network. Using the eigenspectrum and condition number of the critic’s Hessian, we systematically investigate the impact of common architectural design decisions on training dynamics. Our analysis reveals that a novel combination of batch normalization (bn), weight normalization (wn), and a distributional cross-entropy (ce) loss produces condition numbers orders of magnitude smaller than baselines. This combination also naturally bounds gradient norms, a property critical for maintaining a stable effective learning rate under non-stationary targets and bootstrapping. Based on these insights, we introduce xqc: a well-motivated, sample-efficient deep actor-critic algorithm built upon soft actor-critic that embodies these optimization-aware principles. We achieve state-of-the-art sample efficiency across 55 proprioception and 15 vision-based continuous control tasks, all while using significantly fewer parameters than competing methods. Our code is available at danielpalenicek.github.io/projects/xqc.
@@ 1 Introduction
Sample efficiency remains a major challenge in deep rl. Methods that can learn effectively from limited interactions are crucial for applying rl in domains such as robotics (Kober & Peters, 2008; Bohlinger et al., 2025), where generating data on real hardware is costly and time-consuming. Recent advances in model-free rl have primarily been driven by a paradigm of scaling—larger networks, higher update-to-data ratio (utd) ratios, and ever-increasing computational budgets (Nikishin et al., 2022; Nauman et al., 2024; Lee et al., 2025a; Palenicek et al., 2025). These works view architectural improvements primarily as a means to scale up stably and in turn improve sample efficiency. While proven effective, this ‘bigger is better’ paradigm comes at the cost of computational efficiency and overlooks a more fundamental question: Can we improve performance not by adding complexity, but by creating a better-conditioned optimization problem?
To develop this principled understanding, we conduct a systematic investigation into three commonly used architectural components whose roles in rl are often guided by heuristics. First, we examine normalization layers. The rl community frequently uses layer normalisation (ln(Ba et al., 2016), e.g., Ball et al. (2023), in contrast to bn (Ioffe & Szegedy, 2015), which until recently was thought to be problematic in the rl setting (Bhatt et al., 2024). Second, we consider wn (Lyle et al., 2024; Loshchilov et al., 2025; Palenicek et al., 2025) by periodically projecting the network’s weights to the unit sphere, permitted through the normalization layers’ scale invariance property. A technique known to improve the effective learning rate (elr(Van Laarhoven, 2017). Lastly, we study the critic’s loss function. Distributional critics using a ce loss have grown in popularity (Bellemare et al., 2017). The conventional argument for their adoption is that modeling the full distribution of returns provides a better learning signal compared to mean squared error (mse) regression (Farebrother et al., 2024); there is evidence that this loss is easier to optimize (Imani & White, 2018).
@@ 2 Preliminaries
where $Q$ and $V$ are the ‘soft’ parametric critic and value functions, respectively. Minimizing this Bellman error effectively is key to the success of actor-critic methods (Sutton & Barto, 2018). This error is typically minimized with the mean squared regression loss and gradient-based optimization. The distributional c51 algorithm (Bellemare et al., 2017) reformulates the task as a classification problem. Instead of a scalar estimate, the function approximator outputs the logits of a categorical distribution over the full support of $Q$. The Bellman error can then be minimized using a categorical ce loss, which has been shown to improve performance and stability (Farebrother et al., 2024).
@@
${\mathcal{L}}({\bm{\theta}}+\delta{\bm{\theta}})\approx{\mathcal{L}}({\bm{\theta}})+\nabla_{\bm{\theta}}{\mathcal{L}}({\bm{\theta}})\delta{\bm{\theta}}+\frac{1}{2}{\delta{\bm{\theta}}}^{\top}\nabla^{2}_{{\bm{\theta}}}{\mathcal{L}}({\bm{\theta}})\delta{\bm{\theta}},$ (2)
@@ 3.2 Why does cross-entropy outperform the squared error?
For proofs see Appendix J. In practice, we do not require weight decay to attain good performance, and as a result, the Hessian was observed to not always be positive definite. Nonetheless, these results provide a formal intuition for our empirical result that ce losses consistently report smaller condition numbers. Our analysis directly motivates the design of our xqc algorithm, presented next.
@@ 5 Experiments
Figure 6: xqc achieves state-of-the-art sample efficiency across 55 proprioceptive continuous control tasks. We report the iqm auc of normalized returns. Error bars denote 90% sbcis. The right column shows total aggregated performance across the benchmarks (55 tasks). xqc matches or outperforms strong baselines, especially on the hardest DMC and HB tasks, while using a simpler and smaller architecture (see Section 5.3).
@@ 5.1 Sample efficiency results
Reinforcement learning from proprioception. Figure 6 shows that xqc matches our outperforms strong baselines simba-v2, mrq, bro and crossq+wn on all 4 benchmarks. The rightmost column shows that on average xqc performs as well simba-v2 while using significantly less network parameters and a substantially simpler architecture Figure 5. Notably xqc shows exceptional performance on the most complex tasks HB and DMC-hard. These enm induce notoriously difficult and ill-conditioned optimization landscapes. xqc’s superior performance and learning speed suggest that its well-conditioned critic—characterized by a stable elr and bounded gradients as shown in Section 5.2—is fundamentally better equipped to handle the non-stationary targets and bootstrapping errors inherent in these challenging domains.
@@ 5.3 Parameter and compute effciency
Figure 9: xqc is significantly more parameter and compute efficient than competitive baselines. $\sim 4.5\times$ fewer parameters and $\sim 5\times$ fewer flop/s than simba-v2 and bro and $100\times$ fewer flop/s than brc. The computational efficiency translates into a significantly higher training fps measured on the MuJoCo humanoid environment on an nvidia rtx 4090 workstation.
xqc achieves competitive sample efficiency while requiring $\sim 4.5\times$ fewer parameters than simba-v2 (Figure 9). This parameter efficiency directly results in high computational efficiency with $\sim 5\times$ fewer flop/s than simba-v2 and bro and $100\times$ fewer flop/s than brc. We conjecture that xqc’s superior computational efficiency is rooted in its well-conditioned architecture. With respect to practical performance, xqc shows $60\%$ higher frames-per-second (fps) during training, measured on the MuJoCo humanoid on a nvidia rtx 4090 workstation, resulting in significantly faster wall-clock times. xqc’s evaluation performance is measured at 13551355fps, slightly lower the simba-v2 baseline due to its slightly larger policy networks, but remains more than fast enough typical control frequencies seen in deep reinforcement learning (10–50 Hz).
@@ 5.4 XQC scaling behaviour and architecture ablations
[…]
Figure 10 colums 1–3 demonstrate xqc’s scaling ability. xqc improves performance with increasing utd at a similar slope to simba-v2, with a fraction of the parameters (Section 5.3). Similarly, xqc improves or maintains performance with larger and deeper networks. These results demonstrate robustness towards different hyperparameters and xqc’s ability to scale stably, enabled by its well-conditioned architectural design. This property is desirable since increasing compute always represents a trade-off between sample efficiency and wall-clock or energy. It allows practitioners to use their available budget most efficiently. For depth scaling, we also ablate the influence of skip connections but find no significant effect for the investigated depths. Additionally, we investigate the influence of different batch sizes $\in\{128,512\}$ compared to the default of $256$, the use of weight decay, and the impact of action repetition. Neither of these ablations shows a significant influence on xqc. Appendix I presents additional conditioning experiments for the tested components. To demonstrate the synergy between the proposed architectural components, we ablate each of the three main components of the proposed xqc architecture: bn, wn, and the ce distributional critic. Results for all 55 and 5 seeds are shown in Figure 10 with per benchmark ablations in Figure 22. Our analysis confirms that each of the components is vital, especially in the most difficult DMC-hard and HB suits. The largest drop in performance occurs when switching bn for ln. We further investigate ln with half the learning rate, accounting for the increased elr of ln (See Figure 8); however, this leads to even worse performance. This is in line with our finding that ln-based architectures show lower performance due to worse conditioning. While using a mse critic loss has the second most significant influence, the removal of wn is still substantial, but shows the lowest overall impact. In summary, combining each of the three components is vital for xqc’s performance.
@@ 6 Related Work
Our work is positioned at the intersection of several key research areas in deep rl that all attempt to improve sample efficiency. A prevailing trend for improving sample efficiency has been scaling along different axes. Compute scaling, particularly increasing the utd ratio, has recently been a major focus. However, naively increasing utd can lead to instability and overfitting on early experience (Nikishin et al., 2022). Researchers have suggested many different regularizers to stably increase the utd; From full parameter-resets (Nikishin et al., 2022; D’Oro et al., 2022), to critic ensembles (Chen et al., 2021) and drop-out (Hiraoka et al., 2021), to normalization layers like ln (Hiraoka et al., 2021; Lyle et al., 2024; Nauman et al., 2024; Lee et al., 2025a), bn (Bhatt et al., 2024), hyper-spherical normalization (Hussing et al., 2024; Lee et al., 2025b) and spectral normalization (Bjorck et al., 2021). Recently, works have found that the combination of normalization layers together with wn can stabilize the elr, helping against loss of plasticity (Lyle et al., 2024) and also enable scaling rl to high utd ratios (Palenicek et al., 2025). Recent works have worked on developing scaling laws for rl (Rybkin et al., 2025; Fu et al., 2025). Network scaling, is another path authors are exploring to increase sample efficiency. Bhatt et al. (2024) showed that bn allowed them to significantly scale the layer width. Since then, authors have looked into specific ln-based architectures (Nauman et al., 2024; Lee et al., 2025a) and network sparsity (Ma et al., 2025). Lee et al. (2025a) propose a ‘simplicity bias’ score, computed using an FFT and scoring ’simplicity’ higher for functions with lower frequency content across random initializations. This score has no theoretical justification relating it to sample efficiency or generalization. Another line of research attempts to scale network sizes dynamically during training (Liu et al., 2025; Kang et al., 2025). Concurrent work Castanyer et al. (2025) combine second-order optimization and multi-skip residual connections to improve scaling and monitor the trace of the Hessian for deep value-based rl. Orthogonal to our work, recent methods revisit the target network design itself. Vincent et al. (2026) share features between online and target networks to reduce memory while preserving stability, and Hendawy et al. (2026) take the minimum of both estimates for bootstrapping to accelerate learning. Model-based methods scale computation by learning a separate dynamics model, which is leveraged in the rl loop (Janner et al., 2019; Hafner et al., 2020; Palenicek et al., 2023; Hansen et al., 2024).
@@ 7 Conclusion & Future Work
In this work, we shifted the focus from the prevailing pure scaling goal in deep rl and instead focus on improving the critic’s optimization landscape. Through an eigenvalue analysis of the critic’s Hessian, we demonstrate that specific architectural choices, namely batch normalization, weight normalization, and a distributional cross-entropy loss, create a better optimization landscape with a condition number orders of magnitude smaller during learning. This superior conditioning translates directly into learning performance gains. We propose xqc, an algorithm embodying these principles, which achieves state-of-the-art sample efficiency across 70 continuous control tasks from proprioception and vision domains. xqc accomplishes this performance with significantly fewer parameters than competing methods, underscoring that a principled focus on optimization fundamentals can yield greater performance and efficiency than brute-force scaling alone. For future work, insights from xqc can be used to accelerate actor-critic model-based and imitation learning algorithms, e.g., Janner et al. (2019), Watson et al. (2023). xqc performance could be further accelerated by incorporating prior datasets, e.g., Ball et al. (2023).
@@ Acknowledgements
This research was funded by the research cluster “Third Wave of AI”, funded by the excellence program of the Hessian Ministry of Higher Education, Science, Research and the Arts, hessian.AI and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy (EXC-3057/1 “Reasonable Artificial Intelligence”, Project No. 533677015). It was further supported by a UKRI/EPSRC Programme Grant [EP/V000748/1] and partially supported by the German Federal Ministry of Research, Technology and Space (BMFTR) under the Robotics Institute Germany (RIG).
@@ Appendix A Hyperparameters
Table 1 summarizes all proprioception-based experiments’ hyperparameters.
@@
Block design Dense(dim) BN ReLU Dense(dim) ReLU BN Dense(dim) ReLU Dense(4 $\times$ dim) Scaler ReLU Dense(dim) L2 Norm LERP L2 Norm Dense(dim) LN ReLU Dense LN Skip Connection Dense(dim) LN ReLU Dense LN Skip Connection Dense(dim) LN Activation
@@ 10 BRO
Action repeat MuJoCo: 1 DMC: 2 HB: 2 Myo: 2 MuJoCo: 1 DMC: 2 HB: 2 Myo: 2 MuJoCo: 1 DMC: 2 HB: 2 Myo: 2 MuJoCo: 1 DMC: 2 HB: 2 Myo: 2 MuJoCo: 1 DMC: 1 HB: 1 Myo: 1 MuJoCo: 1 DMC: 2 HB: 2 Myo: 2 MuJoCo: 1 DMC: 2 HB: 2 Myo: 2
@@
Block design Dense(dim) BN ReLU Dense(dim) ReLU
@@
Replay buffer capacity 1M 1M
@@
easy: linear(1.0, 0.1, 100K) medium: linear(1.0, 0.1, 500K) hard: linear(1.0, 0.1, 2M) easy: linear(1.0, 0.1, 100K) medium: linear(1.0, 0.1, 500K) hard: linear(1.0, 0.1, 2M)